### v1

Abstract
With the rapid growth of artificial intelligence (AI) in healthcare, there has been a significant increase in the generation and storage of sensitive medical data. This abundance of data, in turn, has propelled the advancement of medical AI technologies. However, concerns about unauthorized data exploitation, such as training commercial AI models, often deter researchers from making their invaluable datasets publicly available. In response to the need to protect this hard-to-collect data while still encouraging medical institutions to share it, one promising solution is to introduce imperceptible noise into the data. This method aims to safeguard the data against unauthorized training by inducing degradation in model generalization. Although existing methods have shown commendable data protection capabilities in general domains, they tend to fall short when applied to biomedical data, mainly due to their failure to account for the sparse nature of medical images. To address this problem, we propose the Sparsity-Aware Local Masking (SALM) method, a novel approach that selectively perturbs significant pixel regions rather than the entire image as previous strategies have done. This simple-yet-effective approach significantly reduces the perturbation search space by concentrating on local regions, thereby improving both the efficiency and effectiveness of data protection for biomedical datasets characterized by sparse features. Besides, we have demonstrated that SALM maintains the essential characteristics of the data, ensuring its clinical utility remains uncompromised. Our extensive experiments across various datasets and model architectures demonstrate that SALM effectively prevents unauthorized training of deep-learning models and outperforms previous state-of-the-art data protection methods.
Keywords:
Unlearnable Example Unauthorized Training Data Protection Adversarial Machine Learning Classification Segmentation.
1 Introduction
The expansion of artificial intelligence (AI) in healthcare has significantly increased the production and storage of sensitive medical data [50, 52]. This data plays a crucial role in advancing research in related fields [68, 32, 52]. The more high-quality, open-source datasets that are available, the more contributions can be made by talented researchers to the development of the field [51, 29, 27, 48, 7, 52]. However, many dataset creators are reluctant to open-source their work due to concerns over unauthorized use, such as training commercial AI models.
Conventional image classification datasets, which are relatively easy to obtain and label by laypersons, as shown by ImageNet’s use of online crowdsourcing [12, 70]. But in the medical field, annotating is a much more complex process requiring specialized knowledge, usually undertaken by experts like radiologists [33, 58, 30]. In addition to identifying the type of disease, detailed annotations regarding the severity of the disease and its relation to surrounding tissues are also necessary. This means that constructing a medical dataset is a labor-intensive task. Unauthorized use could lead to infringement of the creator’s rights, leading to a reluctance to publicly release further data. Moreover, from a patient’s perspective, concerns about their information being exploited for commercial purposes might decrease their willingness to authorize their data for research [34, 61]. Therefore, the construction and release of a medical dataset are not only time-consuming and labor-intensive but also fraught with significant ethical and privacy challenges. The reduction in quality open source datasets, resulting from both of the scenarios mentioned, in turn, slows down the development of medical AI [2, 14, 19].
To defend unauthorized use and encourage sharing data, [25] proposed the technique of contaminating data with imperceptible noise. Models trained on this noise-contaminated data exhibit poor performance for normal utilization. Specifically, these methods create “unlearnable” examples from clean data by adding imperceptible, error-minimizing noise, demonstrating significant protective capabilities. The design of this error-minimizing noise is based on the intuitive idea that an example with a higher training loss could contain more information to learn. Consequently, this noise protects the data by minimizing the corresponding loss, effectively reducing the informativeness of the data.
Nevertheless, applying this method directly to medical images may not be optimal, as it overlooks the unique properties inherent in medical images [42]. The most important property that is overlooked is sparsity [65, 10, 26, 47, 11, 13]. For instance, in cellular microscopy images, even after cropping, there remains a substantial amount of blank background. Similarly, techniques such as CT or tomographic scanning inherently produce sparse data [8, 11, 13]. Previous methods often struggled to pinpoint specific feature regions in medical data, inadvertently emphasizing sparse areas when generating noise. This not only resulted in a significant waste of computational resources but also rendered the noise performance suboptimal, potentially making it ineffective in the workflow of medical data and significantly delaying the open-source release of the data.
To address these challenges, we introduce a novel Sparsity-Aware Local Masking method, which leverages the inherent nature of medical data. Our approach assesses the contribution of pixels to the task based on their gradient and selects pixels with higher contributions for perturbation. This method not only narrows the perturbation search space but also enables the noise generator to focus more on feature regions, yielding noise with stronger protective effects. Additionally, since the protective performance of the noise is entirely derived from feature regions, the performance can be maintained even when cropping large background areas, a common practice in real-world medical workflows. These advancements could significantly motivate medical institutions and researchers to share their data for research or education purposes. In summary, the primary contributions of our research are as follows:
  1. 1.
    We are the first to find that the existing Unlearnable Example overlooks the sparse nature of medical data. Specifically, their performance is not optimal and it is not robust against medical-domain data preprocessing.
  2. 2.
    To address these issues, we propose Sparsity-Aware Local Masking (SALM), which is specifically designed for the medical domain by limiting the perturbation to features improved protection effectiveness and robustness.
  3. 3.
    Experiments on multiple medical datasets from MedMNIST across different modalities, scales, and tasks demonstrate that our SALM achieves SoTA performance and is consistently effective in various medical scenarios. Additionally, we validate the flexible applicability of SALM in real-world situations through case studies.
2 Related Works
2.1 Unlearnable Example
In this paper, we seek to protect data from unauthorized exploitation via a data poisoning approach. Data Poisoning is a technique used to compromise the performance of machine learning models on clean data by deliberately altering training samples. This form of attack has proven effective against both Deep Neural Networks (DNNs) and traditional machine learning methods such as SVM [6]. [45] has highlighted the susceptibility of DNNs to data poisoning, although these attacks typically result in only a modest reduction in DNN performance. However, [62] found there is a clear distinction between poisoned samples and normal samples, which comes at the cost of reduced data usability. Backdoor Attacks represent a specialized form of data poisoning. Traditional backdoor attacks [9, 43] involve the introduction of falsely labeled training samples embedded with covert triggers into the dataset. A relatively new approach within this realm is the creation of Unlearnable Examples [25]. These are considered a more subtle form of backdoor attack, free of labels and triggers. Unlearnable Examples show promising results in protecting data from unauthorized exploitation in various domains and applications [41, 40, 39, 60, 69, 36, 21, 71, 54, 20], such as natural language processing [28], graph learning [39], and contrastive learning [53]. However, the flexibility of protecting medical data has not yet been fully explored. Recently, [42] took a first step to evaluate the performance of conventional image protection methods on medical images. Nevertheless, the methods in [42] are straightforward adaptations from the previous approach and are suboptimal due to a lack of consideration of the intrinsic characteristics of medical data. To better address this, in this study, we leverage the inherent features of medical data to design more effective protection.
2.2 Medical Data Protection
As information technology advances, digital technologies are increasingly integrated into medicine, affecting both clinical treatment and scientific research [55, 51, 29, 27, 48, 7, 52, 68, 32]. However, these developments also present significant risks and challenges regarding patient privacy, medical information breaches occur frequently around the world [3, 31, 49, 46]. Inadequate data protection can lead to substantial harm through the leakage of personal information, such as the disclosure of sensitive health conditions like HIV, potentially leading to social isolation and psychological disorders [18]. Furthermore, incidents of data breaches may also reduce patients’ trust in medical research institutions, making them reluctant to share their data [34, 61]. Earlier studies focused on robust physical protection measures, like using encrypted storage devices [23, 1, 59], establishing firewalls [5, 37], and secure communication transmission modes [17, 35], to safeguard data. [38] introduced a clustering method based on $K$-anonymity algorithm as the building block of privacy-preserving for medical devices’ data. However, as collaboration among various research institutions intensifies and open-source data sharing on the internet becomes more crucial, these methods become less applicable. [4, 66] attempted to use generated data to reduce granularity and thus protect privacy, but this method involves a significant trade-off between information loss and protection efficacy. However, our method remains robust under various conditions during the construction process of real-world datasets and effectively balances high data usability with consistent protection.
3 Methodology
Figure 1: Our SALM method comprises a comprehensive framework that encompasses two primary steps: important pixel acquisition and noise generator training. In the first phase, the model calculates the gradient at each pixel within the image and ranks them, generating a sparse mask through a pre-set $K$ value. In the second phase, the noise generator focuses on perturbing the pixels selected in the previous step and updates its parameters. By implementing this noise, models trained without authorization exhibit poor performance on clean datasets. Conversely, the performance for authorized users remains comparable to that achieved with the original data.
3.1 Problem Statement
Assumptions on Defender’s Capability. We assume that defenders are capable of making arbitrary modifications to the data they seek to protect, under the premise that these modifications do not impair the visual quality. To better simulate real-world conditions, defenders cannot interfere with the training processes of unauthorized users and do not know the specific models they use. Additionally, once the dataset is publicly released, defenders can no longer modify the data.
Objectives. This question is posed in the context of utilizing Deep Neural Networks (DNNs) for the classification of medical images. In a classification task comprising $K$ categories, the clean training dataset and test dataset are denoted as $\mathcal{D}_{c}$ and $\mathcal{D}_{t}$, respectively.
Suppose the clean training dataset consists of $n$ clean examples, that is, $\mathcal{D}_{c}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{n}$ with $\boldsymbol{x}\in\mathcal{X}\subset\mathbb{R}^{d}$ are the inputs and $y\in\mathcal{Y}=\{1,\cdots,K\}$ are the labels and $K$ is the total number of classes. We denote its unlearnable version by $\mathcal{D}_{u}=\left\{\left(\boldsymbol{x}_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{n}$, where $\boldsymbol{x}^{\prime}=\boldsymbol{x}+\boldsymbol{\delta}$ is the unlearnable version of training example $\boldsymbol{x}\in\mathcal{D}_{c}$ ,and $\boldsymbol{\delta}\in\Delta\subset\mathbb{R}^{d}$ is the “invisible” noise that makes $\boldsymbol{x}$ unlearnable. The noise $\boldsymbol{\delta}$ is bounded by $\|\delta\|_{p}\leq\epsilon$ with $\|\cdot\|_{p}$($L_{p}$ norm), and $\epsilon$ is set to be small such that it does not affect the normal utility of the example.
In the specific scenario, a DNN model $f_{\theta}$ parameterized with $\theta$, is trained on $\mathcal{D}_{c}$ to learn the mapping from the input domain to the output domain. For simplicity, we will omit the $\theta$ notation in the rest of this paper. To generate an unlearnable dataset, our objective is to induce the model to learn a spurious correlation between noise and labels: $f_{\theta}:\Delta\rightarrow\mathcal{Y},\Delta\neq\mathcal{X}$, when trained on $\mathcal{D}_{u}$:
[…]
3.2 Error-Minimizing Noise
In an ideal scenario, the generation of noise involves a class-matching process to protect each category within $\mathcal{D}_{c}$. For simplicity, we define the noise generation process on $\mathcal{D}_{c}$ here. Given a clean example $\boldsymbol{x}$, by generating the least disruptive noise $\boldsymbol{\Delta}$ for the training input $\boldsymbol{x}$ by solving the following bi-level optimization problem:
[…]
where $f^{\prime}$ denotes the source model used for noise generation. Note that this is a min-min bi-level optimization problem: the inner minimization is a constrained optimization problem that finds the $L_{p}$-norm bounded noise $\delta$ that minimizes the model’s classification loss, while the outer minimization problem finds the parameters $\theta$ that also minimize the model’s classification loss. However, such an approach does not impose any constraints on the selection of pixels for perturbation. If this method is directly applied to medical data, it overlooks the sparsity inherent in medical datasets.
Input: Training data set $\mathcal{T}$, Training steps $M$, PGD parameters $\alpha_{u}$ and $\rho_{u}$, transformation distribution $T$, the sampling number $J$ for gradient approximation
1 Initialize source model parameter $\theta$.
2 Randomly initialize $\delta^{u}$
3 // Following Eq. (3)
4 for $i$ in $1,\cdots,M$ do
5       Sample minibatch $(x,y)\sim\mathcal{T}$, sample transformation $t_{j}\sim T$
6       Calculate $g_{k}\leftarrow\frac{1}{J}\sum_{j=1}^{J}\frac{\partial}{\partial\delta^{u}}\ell(f^{\prime}_{\theta}(t_{j}(x+\delta^{u})),y)$
7       Determine $k$-th percentile value $g^{0}(k)$ in $g_{k}$
8       for each element $g_{ij}$ in $g_{k}$ do
9             if $g_{ij}\geq g^{0}(k)$ then
10                   Set $M_{ij}=1$
11            else
12                   Set $M_{ij}=0$
13
14      Apply local mask: $\mathcal{P}(g_{k},k)=M\odot g_{k}$  // Following Eq. (5)
15       Update $\delta^{u}\leftarrow\prod_{\|\delta\|\leq\rho_{u}}\left(\delta^{u}-\alpha_{u}\cdot\mathrm{sign}(\mathcal{P}(g_{k},k))\right)$
16       Update source model parameter $\theta$ based on minibatch $(t(x+\delta^{u}),y)$
Output: SALM noise generator $f^{\prime}_{\theta}$, SALM noise $\delta^{u}$
Algorithm 1 Training SALM generator and generating noise.
3.3 Sparsity-Aware Local Masking
To bridge the gap in the biomedical domain, motivated by the observation that the important features in the biomedical image are often sparse as we observed before, we propose a sparsity-aware objective that only modifies a portion of important pixels in the image $\boldsymbol{x}$. Formally, we introduce an additional constraint to limit the noise $\boldsymbol{\delta}$ in terms of $\ell_{0}$ sparsity norm, i.e., $\|\boldsymbol{\delta}\|_{0}\leq m$, we have
$\underset{\theta,\boldsymbol{\delta}:\|\boldsymbol{\delta}\|_{p}\leq\epsilon\text{ and }\|\boldsymbol{\delta}\|_{0}\leq m}{\min}\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}_{c}}\mathcal{L}(f_{\theta}(\boldsymbol{x}+\boldsymbol{\delta}),y).$ (3)
To address the bi-level optimization problem in Eq. (2), existing studies have proposed methods such as iterative generator training [16], target poisoning with a pre-train model [15]. Specifically, in this work, we adopt the iterative generator training framework, with the training termination condition being solely the training steps $M$. When the training step $M$ is sufficient, the noise will be continued optimizing to achieve better performance since there’s no accuracy stop condition like [25]. In each step of noise update, we employ the PGD [44] to solve the constrained minimization problem as follows:
$\delta_{t+1}^{u}=\prod_{\|\delta\|_{p}\leq\rho_{u}}\left(\delta_{t}^{u}-\alpha_{u}\cdot\operatorname{sign}\left(\nabla_{\boldsymbol{x}}\mathcal{L}\left(f_{\theta}^{\prime}\left(x_{t}+\delta_{t}^{u}\right),y\right)\right)\right),$ (4)
where $t$ is the current step in the training process, $\nabla_{\boldsymbol{x}}\mathcal{L}\left(f_{\theta}^{\prime}\left(x_{t}^{\prime}\right)\right)$ is the gradient of loss, $\Pi$ is a projection function that clips the noise back to the refined area around the original example $x$ when it goes beyond, and $\alpha_{u}$ is the step size.
Existing methods do not account for the additional $\ell_{0}$ constraint as proposed in Eq (3). In this paper, we leverage the principle that a pixel’s contribution to the task is proportional to the magnitude of its gradient. Using the gradient as a basis, we rank pixel importance and subsequently generate a localized mask. This approach aims to achieve both high performance and efficiency. Specifically, the gradient corresponding to $\boldsymbol{x}$ is defined as $G_{x}=\nabla_{\boldsymbol{x}}\mathcal{L}\left(f_{\theta}^{\prime}\left(\boldsymbol{x}_{t}^{\prime}\right),y\right)$. The projection $\mathcal{P}$ is then applied to the obtained gradient:
$\mathcal{P}\left(G_{x},k\right)=M\odot G_{x},M_{ij}=\left\{\begin{array}[]{l}1,g_{ij}\geq g^{0}(k)\\ 0,\text{ otherwise }\end{array}\right.,$ (5)
where we seek to modify the top-$k$ percent of the pixels, $g_{ij}$ denotes the value of position $(i,j)$ of the gradient $G_{x}$, and $g^{0}(k)$ is the $k$-th percentile value in $G_{x}$. If $M_{ij}=0$, it means any modification at position $(i,j)$ will result in $\delta_{u}$ not satisfying the constraint conditions of $\|\boldsymbol{\delta}\|_{0}\leq m$.
With projection $\mathcal{P}$ to refine the local region, the perturbation $\delta$ is crafted with a PGD [44] process. Given total PGD steps $K_{a}$, for each iteration $t\in[1,K_{a}]$ the noise is iteratively updated via:
$\delta_{t+1}^{u}=\prod_{\|\delta\|_{p}\leq\rho_{u}}\left(\delta_{t}^{u}-\alpha_{u}\cdot\operatorname{sign}\left(\mathcal{P}\left(g_{k},k\right)_{t},y\right)\right).$ (6)
By purposefully reducing the range of perturbation, this approach enables the noise to cover specific feature regions rather than dispersing throughout the entire color space. This ensures that the noise does not expend effort in sparse regions unnecessarily, leading to better performance and efficiency. The overall framework and procedure are depicted in Fig. 1 and Alg. 1.
4 Experiments
4.1 Setup
Datasets. Our research involved extensive experiments on MedMNIST [63], which is a comprehensive collection of standardized biomedical images spanning 12 distinct datasets. MedMNIST includes the major modalities in medical imaging. This dataset showcases diverse data scales and a range of tasks. Furthermore, to more effectively assess our method’s performance on large, real-world datasets, we also do experiments on the 224x224 datasets from MedMNIST-v2 [64].
Model and Implementation Details. We selected the following well-known models to evaluate our method against benchmark approaches: VGG-11 [57], ResNet-18/50 [22] and DenseNet-121 [24]. For the SALM generator, we select RN-18 as the source model to generate noise. And we set the perturbation noise $\rho_{u}=8/255$ as default. We choose an SGD optimizer for both noise generation and training. Both of their weight decay are set to 5e-4, and momentum is set to 0.9. The initial learning rate is set to 0.1 with a decay rate of 0.1.
Baselines. Since there’s no previous method specifically designed for medical data, we compare our Sparsity-Aware Local Masking (SALM) with the existing SoTA methods in the general image domain. The baseline methods are as following: Target Adversarial Poisoning (TAP) [15], Error-Minimizing noise (EM) [25], Synthesized Perturbation (SP) [67]. None of these three methods impose restrictions on the range of pixels to be perturbed. TAP and EM both target the training loss and craft noise using the gradient information to trigger maximum error or trick models to overfit. SC, on the other hand, imposes hand-crafted linear-separable noise into the data, thereby leading the model to learn only simple noise-label correlation.
4.2 Effectiveness Analysis
Figure 2: Left: The learning curves of ResNet-18 trained on different protected data. Right: The effect of different $K$ values on clean test accuracy(%) for the four datasets.
Table 1: Clean test accuracy (%) of RN-18 trained on datasets across various modalities protected by different methods. The symbol $\downarrow$ in the following context indicates a decrease in accuracy compared to the clean test accuracy.
Dataset Clean AdvT EM SP SALM($K=10$)
PathMNIST 90.7 8.4(normal-$\downarrow$82.3) 13.5($\downarrow$77.2) 12.2($\downarrow$78.5) 11.8($\downarrow$78.9)
ChestMNIST 94.7 38.8($\downarrow$55.9) 78.7($\downarrow$16.0) 63.5($\downarrow$31.2) 27.6(normal-$\downarrow$67.1)
DermaMNIST 73.5 38.7($\downarrow$34.8) 17.1($\downarrow$56.4) 11.5($\downarrow$62.0) 8.7(normal-$\downarrow$64.8)
OCTMNIST 74.3 21.4($\downarrow$52.9) 25.0($\downarrow$49.3) 22.7($\downarrow$51.5) 21.2(normal-$\downarrow$53.1)
PneumoniaMNIST 85.4 66.9($\downarrow$18.5) 65.8($\downarrow$19.6) 70.1($\downarrow$15.3) 63.9(normal-$\downarrow$21.5)
RetinaMNIST 52.4 39.5($\downarrow$12.9) 13.0($\downarrow$39.4) 8.3(normal-$\downarrow$44.1) 12.2($\downarrow$40.2)
BreastMNIST 86.3 44.9(normal-$\downarrow$41.4) 46.2($\downarrow$40.1) 50.0($\downarrow$36.3) 53.2($\downarrow$33.1)
BloodMNIST 95.8 19.6($\downarrow$76.2) 30.5($\downarrow$65.3) 30.5($\downarrow$65.3) 7.1(normal-$\downarrow$88.7)
TissueMNIST 67.6 19.7($\downarrow$47.9) 17.4($\downarrow$50.2) 7.3(normal-$\downarrow$60.3) 21.4($\downarrow$46.2)
OrganAMNIST 93.5 81.0($\downarrow$12.5) 78.1($\downarrow$15.4) 86.1($\downarrow$7.4) 69.0(normal-$\downarrow$24.5)
OrganCMNIST 90.0 70.3($\downarrow$19.7) 72.2($\downarrow$17.8) 74.8($\downarrow$15.2) 64.6(normal-$\downarrow$25.4)
OrganSMNIST 78.2 51.7($\downarrow$26.5) 50.1($\downarrow$28.1) 53.0($\downarrow$25.2) 46.8(normal-$\downarrow$31.4)
[…]
We initially selected PathMNIST for comparison between our method and the conceptually similar error-minimizing and error-maximizing noise. The results in Fig. 2 left show that from the onset of training, both our method and the error-minimizing noise provide effective protection, with our method demonstrating superior performance. Conversely, the efficacy of error-maximizing noise gradually diminished throughout the training. To further evaluate the performance of our method, we directly transferred several conventional image protection methods to medical images and compared them with our approach. The results in Table 1 confirm that our method is not only broadly applicable to different modalities of medical datasets but also outperforms the previous SoTA methods for generic image protection.
4.3 Stability Analysis
Table 2: Clean test accuracy (%) of DNNs trained on the clean training sets ($\mathcal{D}_{c}$) or their unlearnable ones ($\mathcal{D}_{u}$) made by different $K$ value.
[…]
$\mathcal{D}_{c}$
$\mathcal{D}_{u}$
$\mathcal{D}_{c}$
$\mathcal{D}_{u}$
$\mathcal{D}_{c}$
$\mathcal{D}_{u}$
$\mathcal{D}_{c}$
$\mathcal{D}_{u}$
$K=10$ VGG-11
91.7
19.2
73.9
10.9
78.2
21.4
91.7
27.3
RN-18
90.7
11.8
73.5
8.7
74.3
21.2
92.7
25.4
RN-50
94.6
21.9
74.6
9.8
74.3
19.0
94.7
25.4
DN-121
96.4
14.8
75.3
14.5
75.6
13.3
95.1
21.2
$K=30$ VGG-11
90.8
18.3
74.4
11.8
75.5
15.1
92.1
31.4
RN-18
92.9
14.1
73.5
4.0
74.4
12.2
94.6
24.0
RN-50
94.2
17.4
74.4
9.3
77.0
14.3
94.8
30.7
DN-121
94.2
19.7
77.2
11.8
74.9
30.0
94.5
29.9
Different Architectures Selection. Before releasing datasets, we cannot foresee the specific training model unauthorized users might adopt. Hence, ensuring that the protected data remains effective across various models is crucial. While protectors can only select a single source model for noise generation, but unauthorized users are free to choose any model, seemingly placing the former at a disadvantage. However, in reality, the choice of model by users does not impact the efficacy of our method. The results in Table 2 demonstrate that the SALM with the source model ResNet-18 is effective across various models. Furthermore, the protection performance is not necessarily best when unauthorized trained on ResNet-18. This indicates that our method is not limited by the source model and the target model, enabling more effective application in real-world scenarios and facilitating the open sourcing of high-quality datasets. What’s more, the protective effect when $K$ is 30 is not significantly better than when it is 10. This prompts us to delve deeper into the impact of the choice of $K$-value.
Different $K$ Selection. In this section, we delve deeper into our approach to unveil the nuances of our method’s dependency on the choice of $K$ value. $K$ values were selected within a range from 0 to 90 in increments of ten. And we also tried $K=5$ to test the limitation of SALM. When $K$ is set to 0, it implies that the model is trained on clean data. The results in Fig. 2 right show when $K$ is set at 5 or higher, the noise proves to be generally effective. Consequently, in subsequent sections, unless specifically emphasized, we opt for $K=10$ for our experiments. This finding reinforces our assumption that perturbing only a few crucial pixels is enough. What’s more, it highlights the high efficiency of our approach to medical data, which can enable the accelerated release of datasets.
5 Sensibility Analysis
Table 3: Clean test accuracy (%) of RN-18 trained on datasets protected by different methods after three common low-pass filters.
Dataset Filter Clean AdvT EM SP SALM
PathMNIST Mean 91.3 17.9($\downarrow$73.4) 16.4($\downarrow$74.9) 17.1($\downarrow$74.2) 15.6(normal-$\downarrow$75.7)
Median 90.8 22.4($\downarrow$68.4) 7.2(normal-$\downarrow$83.6) 12.7($\downarrow$78.1) 19.3($\downarrow$71.5)
Gaussian 92.1 35.5($\downarrow$56.6) 11.0(normal-$\downarrow$81.1) 18.0($\downarrow$74.1) 17.7($\downarrow$74.4)
DermaMNIST Mean 73.8 45.8($\downarrow$28.0) 31.9($\downarrow$41.9) 23.9($\downarrow$49.9) 21.5(normal-$\downarrow$52.3)
Median 75.1 62.6($\downarrow$12.5) 34.6($\downarrow$40.5) 19.9($\downarrow$55.2) 17.8(normal-$\downarrow$57.3)
Gaussian 74.7 40.1($\downarrow$34.6) 33.1($\downarrow$41.6) 25.8(normal-$\downarrow$48.9) 40.6($\downarrow$34.1)
OctMNIST Mean 77.7 25.7($\downarrow$52.0) 28.3($\downarrow$49.4) 23.8(normal-$\downarrow$53.9) 29.6($\downarrow$48.1)
Median 79.6 23.5($\downarrow$56.1) 25.7($\downarrow$53.9) 23.6($\downarrow$56.0) 22.6(normal-$\downarrow$57.0)
Gaussian 77.7 22.9(normal-$\downarrow$54.8) 26.1($\downarrow$51.6) 23.1($\downarrow$54.6) 29.2($\downarrow$48.5)
[…]
Median 90.9 38.2($\downarrow$52.7) 62.7($\downarrow$28.2) 50.9($\downarrow$40.0) 28.4(normal-$\downarrow$62.5)
Gaussian 91.1 45.1($\downarrow$46.0) 70.8($\downarrow$20.3) 51.4($\downarrow$39.7) 29.8(normal-$\downarrow$61.3)
Resistance to Low-pass Filters. Constrained by limitations in the quality of medical imaging, the acquired images often exhibit noise. Consequently, researchers commonly employ low-pass filters for pre-processing due to their simplicity and efficiency. Therefore, verifying the resilience of our method to such filtering is of paramount importance. We chose three low-pass filters: Mean, Median, and Gaussian, each with a 3$\times$3 window size. The results in Table 3 show that our method retains its effectiveness post-filtering, and the sensibility is lower than other methods, affirming its applicability within actual medical workflows.
Resistance to Cropping. As we consistently highlighted, medical data exhibits greater sparsity compared to conventional data. For instance, when processing microscopic images, physicians or researchers typically reduce the proportion of the background through cropping and magnifying details to facilitate analysis. In this section, we further validate the effectiveness of our SALM after cropping. We calculate the difference in test accuracy on the protected data before and after cropping. To minimize the impact of cropping itself on performance, we add the difference in test accuracy on the clean data before and after cropping, which we denote as the Gap value. The results in Table 5 show after cropping, the performance gap of TAP and EM significantly narrows. This finding confirms that color space disturbances by these methods are random. If perturbations are concentrated in non-feature areas like transparent slides, cropping might result in a loss of protection. Conversely, SALM focuses on pixels that significantly contribute to the model, which ensures that removing the background does not impact the regions with more protective noise and thus preserves the protection.
6 Utility Analysis
Table 4: Test accuracy (%) of RN-18 trained on different kinds of data of PathMNIST.
Table 5: Image similarity scores of different methods.
Method $\mathcal{D}_{c}$ $\text{Cropped}(\mathcal{D}_{c})$ $\mathcal{D}_{u}$ $\text{Cropped}(\mathcal{D}_{u})$ Gap
TAP 91.2 81.9 26.5 12.7 4.5 $\downarrow$
EM 90.6 82.2 17.2 4.7 4.1 $\downarrow$
SP 90.9 81.6 21.3 12.3 0.2 $\uparrow$
SALM 90.9 82.5 14.9 5.3 1.2normal-$\uparrow$
Method SSIM aHash pHash dHash NMI
TAP 94.6 75.1 88.8 77.2 42.7
EM 92.7 77.7 89.1 77.4 42.9
SP 99.9 95.4 99.9 99.3 99.9
SALM 94.4 77.9 89.4 77.2 44.7
Table 5: Image similarity scores of different methods.
Figure 3: The visualization of the origin image and protected image with noise and their corresponding segmentation results.
Medical imaging plays an important role in the diagnosis of a wide range of diseases. Imaging technologies enable doctors to view more detailed structures, thus facilitating more accurate disease diagnoses. Consequently, it is imperative to not only thwart unauthorized training effectively but also to ensure that perturbations in images do not impair the judgment of doctors. Previous studies indicate that when the perturbed radius is smaller than 8, it doesn’t affect visual judgment. However, the implications in the real world of medical data call for an in-depth assessment of image usability, which is the focus of this section.
We start our analysis by evaluating image similarity between the original image and the one protected. Specifically, we utilized the Structural Similarity Index Measure (SSIM), which assesses luminance, contrast, and structure, along with various Hash methods focusing on low-frequency information. The results in Tab 5 reveal that SYN has the highest scores. This is likely due to its structure, consisting of square regions of differing colors with minimal distinctions within these areas. However, compared with other methods, despite being based on feature perturbations, our method does not substantially affect images. This underscores our approach’s ability to effectively preserve image utility.
Furthermore, physicians often use pre-trained models to segment blood vessels or tissues for auxiliary. Therefore, it’s essential that the protection applied to images does not compromise segmentation accuracy. The results in Fig. 3 show that the segmentation results of the protected image are unaffected. To further validate this, we compared the segmentation masks and observed that in 224x224 pixel images, only 0.4% of the pixels showed any difference from the originals. These pixels had inherently low confidence in the original image segmentation, leading to minor perturbations and changes in their results. Overall, the impact of SALM on segmentation performance is negligible, with the protected images still able to guide toward accurate auxiliary results.
7 Case Study: Large-scale Cooperative Dataset
Constructing a large-scale dataset in the real world often requires collaboration among many researchers or institutions. Before the final integration for release, the publisher has no obligation to protect the integrity of the overall dataset. Consequently, it’s crucial for each collaborating entity to implement robust methods to protect their respective data contributions. Previous experiments have assumed access to all data, which is unlikely during the dataset construction phase. Therefore, this section will investigate the efficacy of our approach when only partial data is available, and there is complete unawareness of the remaining data. We set up two scenarios for validation: class-wise combine and sample-wise combine.
Figure 4: The protected framework in the case of the class-combined dataset. The noise generated for the corresponding obtained class from data combined with other support classes still retains its protective effect.
7.1 Class-wise Combined Dataset
We initially conducted a study on the scenario of the class-wise combination. To better explain the experimental setting, we introduce the concept of “support classes,” which is utilized to augment the existing single-class data. We selected two kinds of support classes: medical classes and general classes. The framework is shown in Fig. 4. First, we chose melanoma from the DermaMNIST as a protected target and selected all classes within the PathMNIST as support classes for our experiments. The melanoma data was combined with PathMNIST to create a ten-class dataset while only preserving the SALM noise generated for the melanoma. Subsequently, to broaden the selection range of support classes, we also selected some general image classes (e.g., airplanes) as support classes.
Following the generation of class-specific noise, we utilized the same experimental settings as in previous tests. Using PathMNIST for support classes, the generated noise resulted in a clean test accuracy of 9.9%. Noise generated with general image classes from ImageNet [12] as support classes led to an accuracy of 8.1%. The choice of either type of support class does not affect the noise performance. This demonstrates the flexible applicability of our method.
7.2 Sample-wise Combined Dataset
Table 6: Effectiveness under different unlearnable percentages on PathMNIST with RN-18 model: lower clean accuracy indicates better effectiveness. $\mathcal{D}_{u}+\mathcal{D}_{c}$: a mix of unlearnable and clean data; $\mathcal{D}_{c}$: only the clean proportion of data. Percentage of unlearnable examples: $\frac{\mathcal{D}_{u}}{\mathcal{D}_{c}+\mathcal{D}_{u}}$.
$K$ Value Percentage of unlearnable examples
0% 20% 40% 60% 80% 100%
$\mathcal{D}_{u}+\mathcal{D}_{c}$ $\mathcal{D}_{c}$ $\mathcal{D}_{u}+\mathcal{D}_{c}$ $\mathcal{D}_{c}$ $\mathcal{D}_{u}+\mathcal{D}_{c}$ $\mathcal{D}_{c}$ $\mathcal{D}_{u}+\mathcal{D}_{c}$ $\mathcal{D}_{c}$
$K=10$ 91.2 90.8 91.0 90.0 89.2 89.9 87.2 88.3 89.2 11.8
$K=30$ 90.9 91.3 91.2 90.2 90.0 89.3 90.1 89.2 87.4 18.7
$K=50$ 91.5 89.7 90.9 89.2 90.3 88.5 89.3 86.9 88.7 10.3
Figure 5: Left: The learning curve trained on RN-18 on three different datasets. Right: The t-SNE results of the whole mixed data.
In addition to the scenarios mentioned above, researchers may only be responsible for a portion of the data within the whole dataset. They extensively collect various classes of data and annotate them, then merge them into a complete dataset based on the labels. This spurred our investigation into the effectiveness of randomly selecting a subset of samples for perturbation. Specifically, we applied SALM to a selected percentage of the training data, leaving the remainder untouched and clean. We trained models on this mixed dataset of unlearnable and clean training data, $\mathcal{D}_{u}+\mathcal{D}_{c}$. For comparison, models are also trained on a completely clean dataset, denoted as $\mathcal{D}_{c}$.
The results in Table 6 show a rapid decrease in effectiveness when less than 100% of the data is selected for SALM application, whatever the $K$ is. Surprisingly, applying SALM to as much as 80% of the data results in a negligible effect. Previous studies have shown that both error-minimizing noise and error-maximizing noise have similar limitations in DNNs [25, 56].
To further illustrate this phenomenon, we take 80% as an example and plot the learning curves of ResNet-18 in the following scenarios: 1) Models trained on 20% clean data; 2) Models trained on 80% data processed by SALM; 3) Models trained on a mixture of these two types of data. The results are shown in Figure 5 left. It is observed that data processed by SALM at 80% remains unlearnable for the model, yet the mere 20% of clean data enables the model to achieve excellent performance. The superior performance trained on mixed data arises solely from the clean data contained within. Furthermore, the visualization results in Fig. 5 reveal that the clean data remains mixed in 2D space, whereas the data processed by SALM is almost linearly separable, indicating its unlearnability. Therefore, in the case of sample-wise combined datasets, our method also demonstrates outstanding performance and robust scalability. Overall, our method provides a way for every collaborating entity within large datasets to protect their interests and ensure each data donor’s contribution will not be misused.
8 Conclusion
This work introduces the SALM method, a novel approach dedicated to generating Unlearnable Examples specifically designed for medical datasets. Extensive observation has revealed that medical datasets are inherently sparse, a characteristic not effectively utilized by existing methods for generating Unlearnable Examples. Consequently, the SALM method is designed to selectively perturb only a specific subset of critical pixels. Extensive experimental results demonstrate that the SALM method effectively protects medical images from unauthorized training. Simultaneously, it ensures stability and effectiveness throughout common medical image processing workflows (e.g., filtering and cropping). Additionally, the processed images retain their visual usability, not impeding clinical diagnosis by physicians. Furthermore, we demonstrate that SALM has strong, flexible applicability in practical application scenarios.
Appendix 0.A Datasets Details
Table 7: The details about the datasets we utilized for our experiments. The default size is 28$\times$28, and the corresponding 224 version is only different in size.
Dataset Modality Task Scale
PathMNIST Colon Pathology Multi-Class(9) 107180
ChestMNIST Chest X-Ray Multi-Label (14)&Binary-Class (2) 112,120
DermaMNIST Dermatoscope Multi-Class (7) 10,015
OCTMNIST Retinal OCT Multi-Class (4) 109,309
PneumoniaMNIST Chest X-Ray Binary-Class (2) 5,856
RetinaMNIST Fundus Camera Ordinal Regression (5) 1,600
BreastMNIST Breast Ultrasound Binary-Class (2) 780
BloodMNIST Blood Cell Microscope Multi-Class (8) 17,092
TissueMNIST Kidney Cortex Microscope Multi-Class (8) 236,386
OrganAMNIST Abdominal CT Multi-Class (11) 58,830
OrganCMNIST Abdominal CT Multi-Class (11) 23,583
OrganSMNIST Abdominal CT Multi-Class (11) 25,211
Appendix 0.B More Implementation Details
0.B.1 Hardware Details
The whole experiments are conducted on an Ubuntu 20.04.6 LTS. Experiments For all datasets are conducted on 1 GPU (NVIDIA RTX 4090 24GB) and 16 CPU cores (Intel(R) Xeon(R) Platinum 8352V CPU @ 2.10GHz). Python 3.10.8 and Pytorch 2.1.2 with CUDA 12.1 are used for all the implementations.
0.B.2 Experiments details
PGD. For the PGD process, we set PGD as a random start with total steps $K_{a}=10$, step size $\alpha_{u}=1.6$, and radius $\rho_{u}=8$. Such parameters not only balance protection time and protection performance but also ensure the utility of the data.
Testing. We employ four models introduced in Sec 4.1 for our experiments. All models were trained for 20,000 iterations, and then the accuracy results are reported on their respective data types. For datasets with a resolution of 28x28, the batch size was set to 128, while for datasets of 224x224 resolution, the batch size was reduced to 64. Additionally, for VGG-11, images were padded to a size of 32x32. All other specific parameters were consistent with those used during the noise generation phase.
Appendix 0.C Baseline Details
0.C.1 Target Adversarial Poisoning
This method uncovers the further potential of adversarial examples, originally intended for attacking pre-trained models. Adversarial examples also demonstrate the powerful effects in data poisoning. Specifically, by utilizing the empirical loss maximization problem in place of the bi-level problem according to Eq (7), the poison generation problem is transformed into an adversarial example issue.
$\max_{\delta\in\mathcal{S}}\left[\sum_{\left(x_{i},y_{i}\right)\in\mathcal{T}}\mathcal{L}\left(F\left(x_{i}+\delta_{i};\theta^{*}\right),y_{i}\right)\right]$ (7)
0.C.2 Error Minimizing Noise
We have already introduced EM in the main text. More specifically, we utilize RN-18 as the source model for generation according to Eq. (2). But here we employed the iterative generation framework within REM to achieve enhanced protection effects. The PGD method is employed for solving the inner minimization problem, and the settings are shown in Appendix 0.B.2.
0.C.3 Synthetic Perturbation
This approach induces data to become linearly separable by incorporating multiple square high-frequency patches into images. Specifically, the method initially generates some vertex coordinates in the form of hypercubes in high-dimensional space, then repeats each dimension to introduce local correlations, and finally scales down. For the perturbation radius, we similarly set it to 8/255.
Appendix 0.D More Results
We present additional clean test accuracy results for the 224x224 resolution in Tab 8. The symbol $\downarrow$ in the following context indicates a decrease in accuracy compared to the clean test accuracy.
Table 8: More results of RN-18 trained on 224 version datasets protected by different methods. The symbol $\downarrow$ in the following context indicates a decrease in accuracy compared to the clean test accuracy.
Dataset Clean AdvT EM SP SALM($K=10$)
[…]
Figure 6: The prediction matrix of models trained on data with three processed classes: CAS, CAE and LYM.
Figure 6: The prediction matrix of models trained on data with three processed classes: CAS, CAE and LYM.
Figure 7: Clean test accuracy(%) of DNNs trained on three selected clean and poisoned classes.
0.D.1 Partial-class SALM.
Here, we assume that a complete dataset is already available and can be fully accessed and processed, as an ideal scenario. This serves as the theoretical foundation for our further case studies in real-world situations. In medical datasets, specific classes frequently attract more attention than others. We selected three classes in PathMNIST: CAS, Colorectal Adenocarcinoma Epithelium (CAE), and Lymphocytes (LYM). SALM was applied to all samples within these selected classes. The results in Fig. 7 illustrate that the selected classes become unlearnable. Crucially, as shown in Fig. 7, the misclassification occurred only in the selected classes. These findings underscore the adaptability of our method. Assuming sufficient computational resources and unrestricted access, in a collaborative dataset, researchers can secure the classes they gather. Even in the final combined dataset, these segments retain limited learning utility for their specific categories, thus protecting the interests of contributors.
0.D.2 Case Study: Class-wise Combined Dataset.
In our case study, we solely conducted experiments on the melanoma category. We replicated this procedure for all categories within DermaMNIST and then aggregated the results. For the selection of support classes, we chose PathMNIST, BloodMNIST, and CIFAR-10, randomly selecting five support classes. On the dataset contaminated and subsequently merged in this manner, the test accuracy of the trained RN-18 was merely 10.3%, which is almost equivalent to the protective effect observed when accessing the entire dataset.
Figure 8: More visualization results of the origin image and the corresponding noise and protected image.
Figure 9: More visualization results of the origin image and protected image with noise and their corresponding segmentation results.